OpenStudy (anonymous):

please help solve and graph -10

OpenStudy (cruffo):

\[-10 \leq 4x+2 < 26\] To solve for x, first subtract 2 from each of the three sides. \[-12\leq 4x < 24\] Now divide each of the three sides by 4. \[-3 \leq x <6\]

OpenStudy (cruffo):

the graph will have a shaded dot or a [ on -3 and an open dot or ) on 6, and everything in between will be shaded.

OpenStudy (anonymous):

thank you so much